Session: Feb 2014 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (S. Mariotti at Mon Mar 10 13:02:17 2014) Success:
no known problems

Svetloe (Zborovskiy A. at Mon Mar 10 13:06:05 2014) Success:

Zelenchukskaya (Kiselev P. at Mon Mar 10 13:06:23 2014) Success:

Badary (Zadrutsky at Mon Mar 10 13:06:37 2014) Success:

(Other) (KVN at Mon Mar 10 13:09:42 2014) Success:

Urumqi (Peng Li at Mon Mar 10 13:13:27 2014) Minor failures: RECO HIGH
From NO004 start a normal observation, has been in the snow

Hartebeesthoek (Jonathan Quick at Mon Mar 10 13:17:54 2014) Success:
No known problems

Noto (C. Nocita at Mon Mar 10 13:30:00 2014) Success:RECE
Only left polarization available!

Metsahovi (Minttu Uunila at Mon Mar 10 13:43:17 2014) Minor failures: TSYS
Negative Tsys values, about -900 for all channels

Metsahovi (Minttu Uunila at Mon Mar 10 15:32:39 2014) Failure:RECE
Noticed after experiment that receiver was not locked properly. Now fixed.

Yebes (Daniel Valero at Mon Mar 10 16:14:28 2014) Success:
Not known problems

Torun (K. Borkowski at Mon Mar 10 17:58:55 2014) Minor failures:
Success exept, as usually, two of BBCs giving no or weak fringes.

Shanghai (Bo Xia at Tue Mar 11 06:59:09 2014) Failure:RECE
Receiver power failure(we cooled the system at UT 12:00.)

Onsala (Jun Yang at Tue Mar 11 07:39:08 2014) Success:
No known problems.

Effelsberg (Uwe Bach at Tue Mar 11 08:49:16 2014) Success:

Shanghai (Bo Xia at Wed Mar 12 01:12:28 2014) Success:
Sh: Good fringes to Sh. Clock parameters had to be modified with regard to the a-priori ones from the station.